A pump ejects 12000 kg of water at speed of 4m/s at 40 sec.Find the average rate at which the pump is working.

Grade:12

2 Answers

anjali
7 years ago
Hi..... Rateof working here means power of machine =mgh/t . On putting the values 12000×4×10=480000 . We haven`t use 40sec because it would be same as we have to multiply and divide it which will equal to 1
hero
7 years ago
In my opinion .The pump applied force=m×g=12000×9.8N work done =force ×distance=12000×9.8×40×4Power=rate of work=12000×9.8×4×40÷40=476400 Joule